OpenStudy (anonymous):

Simplify each expression. Use positive exponents. m ^3 n ^–6 p ^0

Nnesha (nnesha):

okay so there is p^0 anything on 0 power is just one

OpenStudy (studygurl14):

Some rules of exponents you will need: \(\large x^{-m}=\frac{1}{x^m}\) \(\large x^0=1\)

OpenStudy (anonymous):

alright

Nnesha (nnesha):

\[x ^{0} = 1\] \[2^{0} = 1\] so what is \[p ^{0} = ???\]

OpenStudy (anonymous):

Hint: \[x^{-n} \implies \frac{ 1 }{ x }\] Any integer/ variable to the power of 0 = 1. For example: \[7^0 = 1~~~\text{and}~~~x^0 = 1\]

Nnesha (nnesha):

and yes like batman said if you have negative power you just have to flip it

OpenStudy (anonymous):

its hard for me to follow im not that good at math

Nnesha (nnesha):

what is p^0 = ????

OpenStudy (anonymous):

1

OpenStudy (studygurl14):

Correct.^

OpenStudy (studygurl14):

Using that, you can simplify... \(\large m^3n^{-6}p^0\) into... \(\large m^3n^{-6}\)

OpenStudy (studygurl14):

What next?

OpenStudy (anonymous):

do we do n^-6 next?

Nnesha (nnesha):

yes right now you have to flip that bec you cannot have negative exponent \[x ^{-3} = \frac{ 1 }{ x ^{3} }\] for example

OpenStudy (anonymous):

how did you loose the negative?

Nnesha (nnesha):

reciprocal x^-3 is same as  so when you flip it sign gonna change if you havae negative ti change to postive and if you positive it change to negative